find the value of w(w+10) given that 2w+5=15
Answer by mananth(16946)

2W+5=15
add -5 to both sides
2W+5-5=15-5
2W=10
divide both sides by 2
2W/2 = 10/2
w=5
..
W(W+10)
plug the value of w
5(5+10)
5*(15)
=75